OpenStudy (lgbasallote):

Prove: \[\huge S_n = \frac{n(a_n + a_1)}2\]

OpenStudy (lgbasallote):

it's an easy proof....but it's still nice to see some professional do it

OpenStudy (lgbasallote):

i wrote that formula wrong...

OpenStudy (vishweshshrimali5):

See \[\S_n = \cfrac{n}{2} (2a + (n-1)d)\]

OpenStudy (lgbasallote):

and..?

OpenStudy (vishweshshrimali5):

Now put 2a +(n-1)d = a + a+(n-1)d = a_1 + a_n

OpenStudy (vishweshshrimali5):

You get ur answer !

OpenStudy (lgbasallote):

i don't think that proves anything...that was formula transformation

OpenStudy (vishweshshrimali5):

see a_n = a+(n-1)d

OpenStudy (vishweshshrimali5):

That really does //////////// :)

OpenStudy (vishweshshrimali5):

You don't want to use the first formula ?

OpenStudy (lgbasallote):

can you prove a_n = a + (n-1)d?

OpenStudy (vishweshshrimali5):

I think yes

mathslover (mathslover):

yep

OpenStudy (lgbasallote):

i think you;re confused....this is proving...not substitution

OpenStudy (shubhamsrg):

suppose we have 1,2,3,4.....100 adding all, we can see it like this (1+100) + (2+99) .... ( 50 + 51) => (101)*50 = (1 + 100) * (100/2) you may generalize this..

OpenStudy (vishweshshrimali5):

And u asked for a proof ;)

OpenStudy (lgbasallote):

that wasn't a proof @vishweshshrimali5 ....it was formula transformation

OpenStudy (lgbasallote):

and substitution isn't proving either.... @shubhamsrg

OpenStudy (vishweshshrimali5):

Whatever........... leave it I can prove that without using a_n formula

OpenStudy (shubhamsrg):

substitution ? o.O

OpenStudy (vishweshshrimali5):

But you will have to accept that a_n = a+(n-1)d That is basic def. of AP

OpenStudy (vishweshshrimali5):

I am also surprised @shubhamsrg O.o

OpenStudy (shubhamsrg):

i am not! ;)

OpenStudy (vishweshshrimali5):

;)

OpenStudy (lgbasallote):

are any of you familiar with mathematical induction?

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (vishweshshrimali5):

You want that proof........

OpenStudy (lgbasallote):

what both of you did was substitution (which is NEVER accepted as proofs)

OpenStudy (vishweshshrimali5):

Let P(n) be the given statement

OpenStudy (lgbasallote):

no. i want other proof

OpenStudy (lgbasallote):

induction is too dull and boring and predictable

OpenStudy (shubhamsrg):

you're amazing you know that ?

OpenStudy (shubhamsrg):

not you ! ;)

OpenStudy (vishweshshrimali5):

@lgbasallote please give us a "hint" of the proof u want

OpenStudy (lgbasallote):

actually...it's surprising @vishweshshrimali5 does know induction...

OpenStudy (shubhamsrg):

hint? lol..

OpenStudy (lgbasallote):

any proof as long as it's valid and not induction

OpenStudy (shubhamsrg):

induction is one of the many great tools we have in mathematics today!

OpenStudy (lgbasallote):

preferrably, i like to see direct proof and contradiction

OpenStudy (lgbasallote):

i like deduction much more

OpenStudy (vishweshshrimali5):

Ok lets try u want to deduce a formula from a continuous pattern

OpenStudy (lgbasallote):

time for the site to go down....

mathslover (mathslover):

bye

OpenStudy (vishweshshrimali5):

bye

OpenStudy (lgbasallote):

so i suppose i'll just ask this again in the fture

OpenStudy (vishweshshrimali5):

will check out later

mathslover (mathslover):

yep